# Uncovering the Functional Roles of Nonlinearity in Memory

### v1

@@ A.4.1 Speech Commands
Figure 12: a: Mean latent trajectories of the first PC for each digit class in the sMNIST task. The trajectories reveal a slow accumulation of evidence along the primary axis, similar to sentiment classification (Fig. 1). Interestingly, distinct branching patterns are visible: the digit ’6’ diverges first, reflecting its characteristic structure in the initial image segments. Digits ’8’ and ’9’ share overlapping paths until around the midway point, aligning with their visual similarity in the upper halves of the images. b: Explained variance ratio of the first $15$ PCs indicates a more complex structure compared to the $2D$ sentiment attractor in Fig. 1.
Figure 14: Analysis of class manifolds for the sMNIST task. (a) Kernel Density Estimates (KDE) of key metrics across different nonlinear configurations ($P=0,50,100$). For the fully linear model ($P=0$), class variance is tightly packed and uniformly distributed, as indicated by low Gini coefficients and Max-Min ratios. Introducing moderate nonlinearity ($P=50$) allows the AL-RNN to strike an effective balance, with a CV close to $1$, signifying well-proportioned variance across class manifolds. This results in distinct spatial separability, as evidenced by lower Gini coefficients and more balanced Max-Min ratios. In contrast, the fully nonlinear model ($P=100$) shows larger Gini coefficients and very wide Max-Min ratios, indicating a much more uneven distribution of variance. This suggests that certain class manifolds (e.g., digit 5) are overly compressed while others are disproportionately spread out, as evidenced in (b): Visualization of the final latent states projected onto the first three PCs for digits 3 (orange) and 5 (green), which represent the most frequent class confusion. At $P=50$, class manifolds are spatially distinct and well-separated, indicating effective utilization of latent space. In contrast, for $P=100$, these manifolds become elongated and entangled, with digit 5 particularly over-compressed. This spatial entanglement directly correlates with the misclassification patterns observed, reflecting an inefficient partitioning of latent space.
Figure 15: Top: First PC of latent activity during the copy task for the best-performing linear model ($65\%$ test accuracy). Bottom: First 3 PCs of freely generated activity.
Figure 16: Distribution of linear subregions during decoding stage for the copy task. (a) The Gini coefficient of bitcode distributions as a function of $P$, reflecting the imbalance in subregion usage. As $P$ increases, the Gini coefficient rapidly rises, indicating that only a limited subset of regions is dominantly occupied. (b) The effective number of linear regions occupied by the AL-RNN, plotted alongside the theoretical maximum (dashed gray line) computed as $\min(2^{P},\text{total samples})$. Despite the potential to explore exponentially many regions, the network only populates a small fraction. (c) Cumulative mass plot for the bitcode distributions sorted by probability, visualized across different values of $P$. Larger $P$ results in most of the probability mass being captured by a small subset of regions.
@@
Figure 17: Distribution of bitcode subregions in the addition task for different levels of nonlinearity $(P)$ for the best-performing models. Across all levels of nonlinearity, we observe a characteristic structure: a small subset of dominant subregions encode the linear integration dynamics, while sparse gating is implemented by a few low-frequency subregions (red). The relative ratio of subregion activity for masked time points is color-coded, and indicates a clear separation of subregions visited during linear integration (blue) and during gating (red), with no region implementing both tasks (see colorbar). While in the more nonlinear settings ($P=10,50$), the linear integration dynamics are distributed over multiple regions, these regions are close neighbors. This could introduce instability compared to the more compact representations seen with sparse nonlinearity, potentially explaining the performance decrease observed with fully nonlinear models. Note that for $P=10\&50$, the y-axis was scaled logarithmically for better readability.

### v2

@@ A.4.2 Classification Tasks
To verify that the trends observed in sMNIST generalize across modalities, we evaluated the AL-RNN on the Speech Commands dataset (Warden, 2018), a ten-class audio classification task using spoken commands sampled at $16$ kHz. We used the same evaluation setup as for sMNIST, with classification based on the final latent state after processing the input sequence. As in the visual domain, we observed that increasing nonlinearity improved classification accuracy up to an intermediate number of nonlinear units, after which performance declined for fully nonlinear models (Fig. 13a). Latent states from models with moderate nonlinearity again exhibited clear class-specific clustering in PC space (Fig. 13b), and bitcode alignment within classes remained high (Fig. 10)—closely mirroring the gating behavior observed in sMNIST (Sect. A.4.2). These results confirm that the latent integration and partitioning mechanisms learned by the AL-RNN generalize robustly across visual and auditory modalities.
A.4.3 Additional Results Addition and Copy Task
@@
Figure 14: a: Distribution of bitcode subregions in the addition task for different levels of nonlinearity $(P)$ for the best-performing models. A small subset of dominant subregions (blue) encode linear integration dynamics, while sparse gating is implemented by a few low-frequency subregions (red). The relative ratio of subregion activity for masked time points shows clear separation between integration and gating functions, with no region implementing both. While more nonlinear settings ($P=10,50$) distribute integration dynamics across multiple neighboring regions, this could introduce instability compared to the compact representations with sparse nonlinearity, potentially explaining performance decreases in fully nonlinear models. Note that for P=10&50P=10\&50, the y-axis is scaled logarithmically. b: Networks often "linearize" certain units, pushing individual activations far from the piecewise-linear boundary such that only a small subset effectively switches. This inductive bias simplifies the optimization landscape by operating in the linear regime where possible, which may explain why fully nonlinear models are less robust. c: Correlation between training and test losses across runs ($r=0.995$). Each point represents a trained model; the dashed line indicates $y=x$. Near-perfect correlation with slightly higher test losses rules out classical overfitting.
Autonomous Dynamics in the Copy Task
To understand how the flow field implements these computations, we further examined the autonomous dynamics of the system. These establish a global k-cycle with a precise period of 100 time steps (exactly half the delay phase duration), as confirmed by a neutral maximum Lyapunov exponent ($\lambda_{max}\approx 0$, Fig. 3b, bottom, see Appx. A.2 for methodological details), for which trajectories neither diverge exponentially nor converge to a fixed point. While the initial transient dynamics are constrained entirely to the lower subregion, the fully developed k-cycle spans both subregions, oscillating between them in a structured manner. This cycle is stabilized through the interplay of two different dominant eigenvalues of the local Jacobians: one with a contracting virtual fixed point ($\lambda_{max}\approx 0.992$) and the other around a diverging fixed point ($\lambda_{max}\approx 1.005$) which drives the initial transient dynamics. Intuitively, once the trajectory has settled onto the orbit that alternates between those two subregions (as visible in Fig. 3b, bottom, where the PWL unit repeatedly crosses the switching boundary), the net effect over a full cycle is neutral expansion $\lambda_{max}\approx 0$), which stabilizes the limit cycle.
Copy Task With Variable Delay
Figure 15: Results for the variable delay copy task. a: Symbol recall accuracy as a function of the number of PWL units P. Performance rises steeply with sparse nonlinearity and peaks at intermediate values of P. Error bars denote standard deviation across seeds. b: Mean activation of individual PWL units aligned to cue onset (dashed red line). During the delay period (steps <0<0), activity remains confined to a largely linear regime. After cue onset, recall requires structured transitions across multiple subregions, reflected in distinct activation patterns across units and time steps.
In the main text, we analyzed the copy task under fixed delays between encoding and recall. We additionally implemented a variable delay version of the task, in which the cue to initiate recall occurred after a randomly chosen interval. In this setting, solutions discovered under fixed delays did not transfer successfully, but required a different strategy for solving the task. However, consistent with the fixed delay task, we find that successful AL-RNNs still maintain latent activity within a single linear subregion throughout the delay period, indicating that storage continues to rely on stable linear dynamics. However, the decoding mechanism is substantially more complex: during the recall phase, trajectories now transition across multiple linear subregions, rather than remaining confined to one. Analysis of the resulting bitcodes shows that symbol identity is reflected in the sequence of subregion switches, with decoding accuracy from the bitcode alone reaching around $60\%$ (compared to a $25\%$ chance level) (Fig. 15b). These results indicate that variable timing shifts the computational burden from storage to decoding. Whereas fixed-delay tasks can be solved by maintaining activity in a single regime and reactivating stored content at a fixed time, the variable-delay setting requires a more flexible, state-dependent decoding process.
Figure 16: Distribution of linear subregions during decoding stage for the copy task. (a) The Gini coefficient of bitcode distributions as a function of $P$, reflecting the imbalance in subregion usage. As $P$ increases, the Gini coefficient rapidly rises, indicating that only a limited subset of regions is dominantly occupied. (b) The effective number of linear regions occupied by the AL-RNN, plotted alongside the theoretical maximum (dashed gray line) computed as $\min(2^{P},\text{total samples})$. Despite the potential to explore exponentially many regions, the network only populates a small fraction. (c) Cumulative mass plot for the bitcode distributions sorted by probability, visualized across different values of $P$.
Figure 17: Training dynamics and performance of AL-RNNs on copy task and addition task. (a) Loss (left) and gradient norm (right) trajectories for models with varying numbers of $P$ PWL units out of 50 (see y-axis labels). Copy task (top): Linear models ($P=0$) show smooth training but fail to reach low loss values. With sparse nonlinearity ($P=1,5$), some models achieve good performance with relatively smooth training, though bifurcations begin to appear. Fully nonlinear models ($P=50$) exhibit highly irregular training with frequent bifurcations and consistently elevated gradient norms. Addition task (bottom): Loss drops faster for models with more nonlinearity, likely since higher model capacity means solutions are discovered more rapidly, followed by fine-tuning. Models with minimal nonlinearity ($P=1$) often fail to find solutions, possibly because fewer nonlinear units reduce the probability of stumbling upon the correct mechanism. Intermediate configurations ($P=10,25$) still show training instabilities where losses suddenly jump up. Fully nonlinear models converge fastest but maintain higher gradient norms throughout. (b) Test accuracy for purely linear networks across latent dimensions $M$ and regularization strengths $\tau$. Linear networks completely fail on the addition task, performing at chance level across all settings. On the copy task, accuracy ranges from $40-70\%$ with apparent random fluctuation across hyperparameters, never approaching optimal performance. This demonstrates that the results from the main text are no overly sensitive to hyperparameter search.
Figure 18: Performance and bitcode distributions on copy task, addition task, and SCAN task. (a) Copy task performance across nonlinearity levels $P$. Best model accuracy (left) peaks at intermediate $P=3$, reaching 100% accuracy. Subregion usage for the best $P=3$ model (middle, $100\%$ accuracy) concentrates in a small number of regions, while the best $P=50$ model (right, $48\%$ accuracy) fragments activity across hundreds of subregions. (b) Addition task test MSE. The best individual model is fully nonlinear ($P=50$, left), while median performance was slightly worse with full nonlinearity. Comparison of the the best $P=50$ model (middle) vs. worst performing models (right) shows that the well-performing model concentrates gating in few subregions with a strongly skewed distributions, while the worst model spreads activity across many subregions. (c) SCAN task accuracy. Best models (left) achieve near-perfect performance across all $P$ values, including $P=128$ (fully nonlinear). However, mean absolute deviation (MAD, middle) increases substantially for fully nonlinear models, indicating higher training variability. Train-test accuracy correlation (right) shows this variance is not due to overfitting but reflects training sensitivity. Fully nonlinear models can still find optimal solutions but do so less reliably.
A.4.4 Nonlinearity Enables Internal Task Switching Through Context-Dependent Routing
We assess the network’s ability to switch between tasks using a context-dependent integration task (Fig. 19a), a design inspired by behavioral paradigms in neuroscience for studying flexible decision-making (Mante et al., 2013). Each trial begins with a one-hot context cue, followed by a sequence of scalar inputs ($T_{seq}=100$). This task requires the network to switch between two internal decision policies based on a contextual cue. Purely linear AL-RNNs cannot solve this task, and plateau at $50\%$ accuracy (Fig. 19b), independent of how many linear units are provided. In contrast, the task is almost perfectly solvable with just one linear and one PWL unit, achieving up to $96\%$ accuracy.
@@ A.4.4 Nonlinearity Enables Internal Task Switching Through Context-Dependent Routing
A.4.5 Nonlinearity Enables Task Switching in a Stimulus Selection Task in Joint Task–Neural Modeling
Figure 21: a: Schematic of the contextual stimulus selection task used in the PFC-1 dataset (Rodgers and DeWeese, 2014). In each trial, animals responded to auditory stimuli based on either the spatial origin (left/right) or the pitch (high/low), depending on the current task rule. b: Latent trajectories from a task-trained 2D model for two identical stimuli under different rules, illustrating context-dependent separation at decision time by the subregion boundary (black horizontal line). c: Task performance across models of increasing nonlinearity (P) and different latent dimensions (M) (left) and misclassification counts by stimulus and block (right). d: Mean firing rate MSE with and without inputs. e: Joint training performance when modeling both task and spike data: test accuracy (left), mean spike rate MSE (middle left), and coefficient of variation (CV) MSE (middle right) and mean spike correlation (right) as a function of the number of nonlinear units. f: Mean activation over time of a representative nonlinear unit that differentiates rules and modulates stimulus-driven activity conditionally.
We next evaluated a structurally similar setting, this time training the AL-RNN not only on the task but also to jointly reconstruct neural population activity from the CRCNS PFC-1 dataset (Rodgers and DeWeese, 2014). This dataset records single-unit activity in rodent auditory cortex during a flexible stimulus-selection task, in which animals alternate between two rules: choosing the correct response based on either the spatial origin or the pitch of the sound (Fig. 21a). The two stimulus dimensions are presented simultaneously but independently, with either a low or high warble (pitch cue) played on both sides, and a pure tone (side cue) played either on the left or right. Depending on the current rule, the animal must attend to either the pitch (low = "go", high = "no-go") or the side (left = "go", right = "no-go") of the auditory input. This setup creates contextual ambiguity: for example, an identical stimulus combination of a high-pitched warble and a tone on the left requires different responses ("go" under the spatial rule, "no-go" under the pitch rule). Hence the task studies linear and nonlinear neural computations in context-dependent behavior. Preprocessing steps and selection criteria are detailed in the Appx. A.3. To ensure a consistent task representation, we restricted training and evaluation to the $723$ trials in which the animal made the correct choice.
Task-training the AL-RNN
We first trained an AL-RNN to solve the task based purely on structured temporal input, independent of neural data. Each trial mimicked one of the 723 empirically observed trials, beginning with a brief context cue indicating the active rule (spatial or pitch), presented for 100 ms (two time steps), followed by a one-second delay. Then, an auditory stimulus was presented (one-hot encoded for pitch and side), and the model was required to produce a binary decision (go/no-go), which was read out linearly from the latent state at the empirically observed decision time $\bm{z}_{t_{dec}}$ (see Appendix A.3). Despite the task’s contextual nature, we find that a 2D model is already often sufficient to linearly separate identical stimuli based on their preceding rule cue, and linear 3D models often solved the task perfectly. Systematic evaluation confirmed that increasing capacity beyond this dimension does not yield substantial gains in task accuracy alone (Fig. 21c, left), and that errors were largely restricted to contextually ambiguous stimuli (e.g., Right–Low under the spatial rule; Fig. 21c, right).
Task Information Improves Spike Reconstruction
We next used the same AL-RNN to jointly model both the animal’s behavior and the neural spike trains. That is, the network received identical inputs as in the task-only training—context cue, delay, and stimulus—and was required not only to produce the correct decision at the appropriate time, but also to reproduce the observed spiking activity across neurons from the same latent states of the AL-RNN (Eq. 1). To achieve this, we trained the AL-RNN within the Multimodal Teacher Forcing (MTF) framework (Brenner et al., 2024b), which allows direct training on non-Gaussian spike data via a Poisson decoder. The decoder is hierarchically conditioned on a 5-dimensional, trial-specific feature vector (Brenner et al., 2025), enabling it to capture slow drift in spike statistics while preserving a shared task representation in the AL-RNN (see Appx. A.1.4 for training details and Fig. 23 for an analysis of the learned features).
We found that including task information markedly improved spike reconstruction (Fig. 21d), suggesting a strong alignment between latent decision dynamics and observed neural variability. Quantitatively, spike rates and coefficients of variation were well captured (see Fig. 22a). To evaluate the model’s temporal reconstruction, we correlated recorded and generated spike trains, yielding a mean correlation of $r=0.36\pm 0.002$. This approached the correlation between separate Poisson samples drawn from the same latent trajectory ($r=0.371\pm 0.057$), which serves as an approximate upper bound given the high intrinsic variability of sparse cortical spiking.
Nonlinearity Supports Integration In Joint Modeling
In the joint setting, nonlinearity began to play a more significant role. Increasing the number of nonlinear units improves both task accuracy and spike reconstruction ($r\approx-0.37,p=0.005$, low firing rate MSE/higher task accuracy are better, Fig. 21e). In contrast, linear models plateau around $87\%$ accuracy and exhibit an almost identical error pattern: they consistently respond “go” to all trials with left or low cues, regardless of the active rule, thus failing to resolve the contextual ambiguity (see Fig. 21c, right). In many such nonlinear models, we observed the emergence of single units that effectively gated the task rule, becoming active in one rule context and inactive in the other, with stimulus integration occurring only within the active subregion (Fig. 21f). Quantitatively, we examined the linear subregion active at the time the input is first presented (after one second) and compared which subregions were active in blocks 2 and 4 (the two conditions where both spatial and pitch cues are present but different rules are active; see Appx. A.3). In successful models ($99\%$ accuracy), trials from these blocks were routed into almost entirely distinct subregions (on average, $97\%$ of trials fell into block-specific linear subregions), while in unsuccessful models this separation was much less pronounced (only $32\%$). Task accuracy was strongly correlated with this separation score (ratio of trials falling into non-overlapping subregions for different rules, $r=0.66$). This shows a clear link between successful models leveraging nonlinearity to assign different task rules to distinct linear regions, allowing them to react differently to identical stimuli depending on context. This mechanism is illustrated for a successful task-trained 2D model in Fig. 21b.
@@ A.4.5 Nonlinearity Enables Task Switching in a Stimulus Selection Task in Joint Task–Neural Modeling
Figure 24: Bitcode usage from AL-RNN training on a multi-task paradigm. a: Left: Bitcode similarity matrices computed from Jensen-Shannon divergence between task-specific subregion distributions, as in Fig. 5. Right: Task-specific bitcode usage distributions corresponding to each $P$ level. For $P=1$, only 2 possible bitcodes exist (representing the two sides of a single PWL boundary), and tasks cleanly partition between them, e.g., DelayPro and DelayAnti use opposite subregions, consistent with their opposite but individually linear task requirements (see also Fig. 25 for example latent trajectories). At $P=5$, with $2^{5}=32$ possible bitcodes, distributions broaden but remain concentrated on small subsets, with some tasks showing distinct peaks (e.g., Match2Sample, CtxIntMod1/2) while others overlap. At $P=64$, with $2^{64}$ possible bitcodes, distributions become extremely sparse and diffuse across the top 200 most-used regions, and tasks show almost no shared structure. b: Consistency of learned representations across runs with 20 samples per task via correlation of task similarity matrices (from a) across 5 independent training runs. Sparse nonlinearity (P=416P=4-16) produces highly robust subregion assignments. Representations also remain similar across adjacent sparsity levels (e.g., $P=4$ to $P=8$ shows correlations $\sim 0.5-0.7$), indicating that in the low data limit, the shared nonlinear structure is robustly captured.
Figure 25: Latent trajectories of an AL-RNN with $P=1$ trained on the multi-task suite for the pro and anti versions of the delay task, and the two modalities in the contextual integration task. The network leverages the single PWL boundary (black horizontal line) to separate the task mechanism for the two different task versions.
@@
Figure 27: Comparison of three architectural variants for both Linear Transformer (Katharopoulos et al., 2020) and S4 (Gu et al., 2022) models: Fully Linear models contain no nonlinear activations anywhere in the architecture, Standard models include nonlinearity in every layer (ReLU in feedforward networks for Transformers, GELU in every S4D block for S4); and Sparse Nonlinear models only use nonlinearity at the final layer after all attention or state-space blocks, keeping intermediate computations linear. Note that the optimization is not directly comparable to the training of the AL-RNN. While we implemented causal versions of these architectures to maintain meaningful task structure, even with causal masking, these models fundamentally differ from first-order Markovian RNNs, with their multi-layer structure provides a view of the sequence that enables linear operations to solve tasks like addition (which our proof shows is impossible for purely linear Markovian models). (a) Linear Transformer performance. On the addition task (left), standard Transformers with distributed nonlinearity converge slower and converge at higher error than both fully linear and sparse nonlinear variants, which perform comparably. Test performance closely mirrors training, with standard Transformers reach $\sim$10-3 MSE while others achieving near $10^{-4}$. On the copy task (right), sparse nonlinear models substantially outperform both alternatives, reaching near-perfect accuracy (>>95%) on both train and test sets. Standard Transformers show intermediate performance, while fully linear models converge much slower and exhibit high variance. (b) S4 model performance. For the addition task (left), all variants converge to similar final performance ($\sim$10-3 MSE) with, with the nonlinear variants performing slightly better. On the copy task (right), all three variants achieve near-perfect performance, though sparse nonlinear models converge slightly faster than the alternatives. Overall, the architecture-dependent benefits of sparse nonlinearity observed in RNNs partially transfer to attention-based models but are less pronounced in state-space models for the two considered tasks.